### v1

# Temperature of McVittie Black Holes in Cavities

Abstract
In this paper, we study the temperature of the McVittie black holes in cavities. The cavities are isothermal boxes that can make the McVittie black holes thermodynamically stable. The McVittie black holes are dynamical black holes, so the traditional action method can hardly be used to calculate the cavity temperature. We find that the cavity temperature $T_{c}$ can be derived from the McVittie black hole temperature $T_{+}$ and the Tolman relation. By setting $H^{2}$ to be $\Lambda/3$, we also get the cavity temperature of de Sitter black holes, which are interestingly the same with the results from the action method, but our method is much more easier.
I Introduction
In previous workAbdusattar:2022bpg , we investigated the thermodynamics of the McVittie black hole. Found in 1933, the McVittie black holeMcVittie:1933zz ; Faraoni:2007es ; Kaloper:2010ec ; Nandra:2011ug ; Faraoni:2012gz can be regarded as a black hole embedded in the spatially flat Friedmann-Robertson-Walker universe, which is similar to the de Sitter black hole. Like the de Sitter black hole, McVittie black hole has two horizons, the black hole horizon and the cosmological horizon, both of which can evaporate with different temperature. One can treat the two horizons as two separate thermodynamic systems, establish the first law of thermodynamics and so on for the two horizons respectively. However, the two temperatures are not equal, so the McVittie black hole is not in the equilibrium state.
In order to make the McVittie black hole thermodynamically stable, one can place a cavity between the black hole horizon and the cosmological horizon as is often done for de Sitter black holesSimovic:2018tdy ; Marks:2021fpe . It has a fixed temperature and allow the black hole to be in the equilibrium state. In this paper, we use this method to stabilize the McVittie black hole and study its temperature. Unlike the de Sitter black hole, we can not integrate the action of the McVittie black hole, because one can not find a static patch or coordinate system for McVittie black hole. Therefore, we can not get the temperature at the horizons from the action. Fortunately, we have McVittie black hole temperature that can further give us the cavity temperature. The McVittie black hole temperature $T_{+}$ is defined from the surface gravity at the black hole horizon. Start from the $T_{+}$, we can get the cavity temperature $T_{c}$ by using the Tolman relation. In flat spacetime, equilibrium means temperature is the same in every position, but it is not the case for curved spacetime. It is found that in curved spacetime or a gravitational system, temperature can be different even equilibrium is established. For a black hole, the temperature of the Hawking radiation cools down as it travels away from the horizon due to the redshift effect. It has an extreme high temperature near the horizon and is called the Hawking temperature if measured at an infinite far place. Temperature at different distance satisfies the Tolman(or Tolman-Ehrenfest) relationRovelli:2010mv , which says that the multiplication of temperature and the norm of the Killing vector field is a constant, which is actually equivalent to the redshift formula. In this paper, we use this relation to calculate the temperature at $r_{c}$ where the cavity is placed. If the Hubble parameter is treated as a constant, i.e. $H^{2}=\Lambda/3$, the McVittie black hole becomes the de Sitter black hole, so from the result we can also get the cavity temperature of the de Sitter black hole. Interestingly, our result is the same with the result from the previous action method. We further investigate the charged McVittie black holeGao:2004cr ; Faraoni:2014nba ; Guariento:2019ock and get the cavity temperature. Once again, from the result, we get the same expression for the charged de Sitter black hole. These could be regarded as consistency checks, but our method is much more easier.
The paper is organized in the following way. In Sec.II, we introduce the McVittie black hole and calculate its temperature at the cavity. In Sec.III, we introduce the charged McVittie black hole and calculate its temperature at the cavity. In Sec.IV, we give the conclusions.
II Cavity Temperature of the McVittie Black Hole
The metric of the McVittie black hole can be written asNandra:2011ug
$\mathrm{d}s^{2}=-\left(1-\frac{2m}{r}-H^{2}r^{2}\right)\mathrm{d}t^{2}-\frac{2Hr}{\sqrt{1-\frac{2m}{r}}}\mathrm{d}t\mathrm{d}r+\frac{1}{1-\frac{2m}{r}}\mathrm{d}r^{2}+r^{2}(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\varphi^{2}),$ (1)
where $H\equiv\dot{a}(t)/a(t)$ is the Hubble parameter and we use different signature with that paper. The apparent horizon of a dynamical spacetime satisfies
$\nabla^{a}r\nabla_{a}r=0,$ (2)
where $a=0,1$ with $x^{0}=t,x^{1}=r$, so for the McVittie black hole (1), this condition results to
$1-\frac{2m}{r}-H^{2}r^{2}=0.$ (3)
It has two solutions generally, and we denote the smaller one as $r_{+}$, i.e. the black hole horizon, and the larger one as $r_{cosmo}$, i.e. the cosmological horizon. From the above condition, one can get a useful relation
$\dot{r}_{+}=\frac{2H\dot{H}r_{+}^{3}}{1-3H^{2}r_{+}^{2}}.$ (4)
The surface gravity at radius $r$ is defined asAbdusattar:2022bpg
$\kappa:=\frac{1}{2\sqrt{-h}}\frac{\partial}{\partial x^{a}}\left(\sqrt{-h}~h^{ab}\frac{\partial r}{\partial x^{b}}\right),$ (5)
where $a,b=0,1$ $h^{ab}$ is the metric at the $(t,r)$ space and $h$ is its determinant, and inserting the metric (1), one can get
$\kappa=\frac{m}{r^{2}}-\frac{\dot{H}r}{2\sqrt{1-\frac{2m}{r}}}-H^{2}r.$ (6)
At the black hole horizon $r_{+}$, the surface gravity can be written as
$\kappa_{+}=\frac{1}{2r_{+}}\left(1-3H^{2}r_{+}^{2}\right)\left(1-\frac{\dot{r}_{+}}{2H^{2}r_{+}^{2}}\right),$ (7)
so one can get the Hawking temperature of the black hole horizon
$T_{+}=\frac{1}{4\pi r_{+}}(1-3H^{2}r_{+}^{2})\left(1-\frac{\dot{r}_{+}}{2H^{2}r_{+}^{2}}\right),$ (8)
where (3) is used.
At the cavity with radius $r_{+}<r_{c}<r_{cosmo}$, the temperature $T_{c}$ is fixed with the temperature of the Hawking radiation there, which is different with $T_{+}$ but can be obtained from the Tolman relation [^1] The time-like observer is $\xi^{a}=(\partial/\partial t)^{a}$, so its norm is $\xi||=\sqrt{|g_{ab}\xi^{a}\xi^{b}|}=\sqrt{|g_{00}\xi^{0}\xi^{0}|}=\sqrt{|g_{00}|}=\sqrt{-g_{00}}$.
[…]
The McVittie black hole (1) can reduce to de Sitter black hole by fixing $H^{2}=\Lambda/3$, and in this case we have $\dot{r}_{+}=0$ from (4). Therefore, from the above result one can also get the cavity temperature of de Sitter black hole,
$T_{c}=\frac{1-\Lambda r_{+}^{2}}{4\pi r_{+}\sqrt{\left(1-\frac{r_{+}}{r_{c}}\right)\left[1-\frac{\Lambda}{3}(r_{+}^{2}+r_{+}r_{c}+r_{c}^{2})\right]}},$ (10)
which is exactly the same with the result by the action methodSimovic:2018tdy .
III Cavity Temperature of the Charged McVittie Black Hole
Similar to the charged de Sitter black hole, the solution of the charged McVittie or cosmological black hole has also been found. The metric of the charged McVittie black hole can be written asGuariento:2019ock
$\mathrm{d}s^{2}=-\left(N^{2}-H^{2}r^{2}\right)\mathrm{d}t^{2}-\frac{2Hr}{N}\mathrm{d}t\mathrm{d}r+\frac{1}{N^{2}}\mathrm{d}r^{2}+r^{2}(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\varphi^{2}),$ (11)
where $N\equiv\sqrt{1-2m/r+Q^{2}/r^{2}},H\equiv\dot{a}(t)/a(t)$, and $Q$ is the charge.
The apparent horizon of the charged McVittie black hole also satisfies (2), which in this case can be written as
$1-\frac{2m}{r}+\frac{Q^{2}}{r^{2}}-H^{2}r^{2}=0,$ (12)
and the smaller solution or the black hole horizon is also denoted by $r_{+}$. From the above condition, we can also get
$\dot{H}=-(r_{+}^{2}-3mr_{+}+2Q^{2})\frac{\dot{r}_{+}}{Hr_{+}^{5}}.$ (13)
Insert the metric (11) into the definition (5) of the surface gravity, one gets the surface gravity at radius $r$ of the charged McVittie black hole,
$\kappa=\frac{m}{r^{2}}-\frac{Q^{2}}{r^{3}}-\frac{\dot{H}r}{2N}-H^{2}r.$ (14)
At the black hole horizon, the surface gravity can be written as by using (3)
$\kappa_{+}=\frac{1}{2r_{+}}\left(1-3H^{2}r_{+}^{2}-\frac{Q^{2}}{r_{+}^{2}}\right)\left(1-\frac{\dot{r}_{+}}{2H^{2}r_{+}^{2}}\right),$ (15)
so the Hawking temperature can be obtained
$T=\frac{\kappa_{+}}{2\pi}=\frac{1}{4\pi r_{+}}\left(1-3H^{2}r_{+}^{2}-\frac{Q^{2}}{r_{+}^{2}}\right)\left(1-\frac{\dot{r}_{+}}{2H^{2}r_{+}^{2}}\right).$ (16)
Therefore, from the Tolman relation the temperature at the cavity $r_{c}$ of the charged McVittie black hole is
[…]
If one sets $H^{2}=\Lambda/3$, the charged McVittie black hole reduces to the charged de Sitter black hole, so we can get the cavity temperature of the charged de Sitter black hole
$T_{c}=\frac{1-\Lambda r_{+}^{2}-\frac{Q^{2}}{r_{+}^{2}}}{4\pi r_{+}\sqrt{(1-\frac{r_{+}}{r_{c}})\left[1-\frac{Q^{2}}{r_{+}r_{c}}-\frac{\Lambda}{3}(r_{c}^{2}+r_{+}^{2}+r_{c}r_{+})\right]}},$ (18)
which is still equivalent to the result from the action methodSimovic:2018tdy .
IV Conclusions
In this paper, we investigate the temperature of the McVittie black hole and the charged McVittie black hole in a cavity with radius $r_{c}$. The cavity temperature is derived from the black hole Hawking temperature and the Tolman relation. We also get the cavity temperature of the de Sitter black hole and the charged de Sitter black hole respectively by setting $H^{2}=\Lambda/3$, which are interestingly the same with the results from the action method. The method in our work is much easier in computations and can be used to similar investigations.
Acknowledgment
This work is supported by National Natural Science Foundation of China (NSFC) under No.12465011 and East China University of Technology (ECUT) under DHBK2023002.